OpenStudy (anonymous):

prove this statement through mathematical induction: (3^n) - 1 is divisible by 2.

OpenStudy (anonymous):

There is at least one \(n\) such that\[\frac{3^n-1}{2}\in\mathbb Z\]We can test this by picking any positigve integer. Does \(\frac{3^{n+1}-1}{2}\in\mathbb Z\) also?\[\frac{3^{n+1}-1}{2}=\frac{3^n3-1}{2}=\frac{3}{2}3^n-\frac{1}{2},\left\{\frac{3^n}{2}=a+\frac{1}{2}\mid\forall n\because3^n\in\mathbb Z,\frac{3^n}{2}\notin\mathbb Z\right\}\]I think you can do it from here?

OpenStudy (anonymous):

\(3^n-1\) is even. Why bother using induction?

OpenStudy (anonymous):

Because sometimes when you do the same thing multiple times you get a good grade.

OpenStudy (anonymous):

I see what you're doing, but it needs to be done using induction :/

OpenStudy (callisto):

This is the only type of question I love :P For n=1 (3^n) - 1 = (3^1) -1 =2, which is divisible by 2 So, the proposition is true for n=1 Assume (3^k) - 1 is divisible by 2 for some positive integers k, that is (3^k) - 1 =2m, where m is a positive integer. Then n=k+1 (3^(k+1)) -1 = 3 x (3^k) -1 = 3( 2m +1) -1 = 6m +3 -1 = 6m +2 = 2(3m+1), where 3m+1 is a positive integer therefore (3^(k+1)) -1 is divisible by 9 therefore the proposition is also true for n=k+1 By principle of MI, the proposition is true for all positive integers n

OpenStudy (callisto):

If the question asks you to do so, it's difficult to ask why :S

OpenStudy (anonymous):

thanks! but I don't get this line: = 3 x (3^k) -1

OpenStudy (anonymous):

That line may have looked like this:\[3^1*3^k-1=3*3^k-1\]

OpenStudy (callisto):

Thanks for the explanation @rchstudy I didn't realise that there was a reply. Sorry!!

OpenStudy (anonymous):

no prob @Callisto